the measure of an angles supplement is 44 less than the measure of the angle find the measure of the angle and its supplement i am so confused and lost please help thank you so much
Answer by stanbon(57282)

the measure of an angles' supplement is 44 less than the measure of the angle find the measure of the angle and its supplement
------
Let the angle be "x".
Then the supplement is "180-x".
--------------------------------
Equation:
supplement = angle - 44
180-x = x - 44
2x = 224
x = 112 degrees (the angle)
180-x = 68 degrees (the supplement)
======================================
Cheers,
Stan H.
=====================